A football quarterback shows off his skill by throwing a pass 45.20m downfield and into a bucket. The quarterback consistently launches the ball at 38.00degrees above horizontal, and the bucket is placed at the same level from which the ball is thrown.

QA) What initial speed is needed so that the ball lands in the bucket?

QB) By how much would the launch speed have to increase if the bucket is moved to 47.80m downfield?
1
Expert's answer
2020-02-17T09:27:07-0500

A)

"R=\frac{v^2\sin{2(38)}}{g}"

"45.2=\frac{v^2\sin{76}}{9.8}"

"v=21.37\frac{m}{s}"

B)


"47.8=\frac{v'^2\sin{76}}{9.8}"

"v'=21.97\frac{m}{s}"

"v'=21.97-21.37=0.60\frac{m}{s}"